# SCINE — Software for Chemical Interaction Networks

### v1

@@ II.5.1 Field of application
If a system of interest is too large for a full quantum mechanical description, a quantum–classical hybrid model must be constructed. Such models generally describe a small subsystem quantum mechanically (QM) and the remainder with a molecular mechanical (MM) model which parametrizes inter- and intramolecular interactions based on experimental or calculated reference data. The module SCINE Swoose provides such MM models, the framework to combine its models with any existing electronic structure theory to construct a QM/MM hybrid model, and workflows specialized for nanoscopic systems. Generally, MM models can be difficult to construct due to the structural complexity of nanoscopic systems and the necessity of MM models to define bonds and charges a priori. Swoose overcomes this hurdle by providing a general workflow for structure preparationCsizi and Reiher (2023) and a system-focused atomistic model (SFAM)Brunken and Reiher (2020) that can be generated for each system of interest with automated QM reference data generation by fragmenting the system, carrying out numerous structure optimizations in parallel, and fitting force constants with a partial Hessian procedure. The generated structural and MM model can then be leveraged easily in automated proceduresCsizi, Steiner, and Reiher (2023). Reaction mechanisms can be studied with hybrid models, for which the QM region can be automatically determined by another algorithm implemented in Swoose.Brunken and Reiher (2021) It generates multiple candidates for the QM region around an atom alongside multiple large reference systems, calculates the first-order partial derivatives for each nucleus, i.e., the atomic forces, and compares them against those in the reference systems. The algorithm assumes that small errors in atomic forces are a good descriptor together with energies as shown in Ref. 16. It then selects the smallest QM system that features errors below a given threshold. Compared to other approaches to determine a QM regionSumner, Söderhjelm, and Ryde (2013); Liao and Thiel (2013); Kulik et al. (2016); Karelina and Kulik (2017); Brandt and Jacob (2023a, b), to determine protonation statesDolinsky et al. (2004, 2007); O’Boyle et al. (2011); Olsson et al. (2011); Anandakrishnan, Aguilar, and Onufriev (2012); Sastry et al. (2013); Riojas and Wilson (2014); Bochevarov et al. (2016); Eastman et al. (2017); Reis et al. (2020), and to fit a system specific modelGrimme (2014); Vanduyfhuys et al. (2015); Welsh and Allison (2019); Behler and Csányi (2021); Friederich et al. (2021); Unke et al. (2021); Deringer et al. (2021); Musil et al. (2021), Swoose exploits the straightforward parallelization of multiple calculations within the SCINE framework and provides a complete pipeline from experimental data to systematic reaction explorations with QM/MM hybrid models, including real-time calculationsCsizi, Steiner, and Reiher (2023).
II.5.2 I/O
@@
The algorithms for structure preparation, model generation, and quantum region selection can be carried out with Swoose with an input file following the YAML syntax, which allows one to specify all required settings in a straightforward way. Additionally, Swoose can carry out single-point energy calculations directly, structure optimizations, and molecular dynamics simulations on the command line.
II.5.3 API
@@ II.7.1 Field of application
Highly accurate electronic structure calculations require one to take electron correlation into consideration. Even though closed-shell molecules can be routinely and reliable calculated black-box coupled-cluster modelsBartlett (2024), strongly correlated systems require an active orbital space for a multi-configurational description with subsequent treatment of the dynamic correlation. Since the manual selection of an active space requires chemical intuition and experience, this step is often time consuming, difficult to reproduce, and hardly possible for non-prototypical complex system such as the one in Ref. 132. Especially for methods which can handle large active spaces of more than 100 orbitals, like DMRGWhite (1992, 1993); Baiardi and Reiher (2020) and full configuration interaction quantum Monte CarloBooth, Thom, and Alavi (2009); Cleland, Booth, and Alavi (2010), a manual selection is neither feasible nor reliable. Although many approachesPulay and Hamilton (1988); Bofill and Pulay (1989); Tishchenko, Zheng, and Truhlar (2008); Bao et al. (2016, 2017); Sayfutyarova et al. (2017); Bao et al. (2018); Khedkar and Roemelt (2019); Sayfutyarova and Hammes-Schiffer (2019); Jeong et al. (2020); Li, Gagliardi, and Truhlar (2020); Gieseking (2021); Golub et al. (2021); Khedkar and Roemelt (2021); Lei, Suo, and Liu (2021); Levine et al. (2021); Oakley, Gagliardi, and Truhlar (2021); King and Gagliardi (2021); Weser et al. (2022); Casanova (2022); Cheng, Xie, and Ma (2022); King et al. (2022); Kaufold et al. (2023); Golub et al. (2023), have been devised to address this problem (at least in parts), we have demonstrate in 2016 a viable, reliable, and truly fully automated selection approach based on quantum information measures, the autoCAS approach Stein, von Burg, and Reiher (2016); Stein and Reiher (2016); Stein2017; Stein and Reiher (2019); Bensberg and Reiher (2023b), a development, which has not been paralleled so far.
@@
The autoCAS algorithmStein and Reiher (2016, 2019) exploits the fact that DMRG can converge a qualitatively correct wave function with less effort than what would be required for a converged energy.Moritz and Reiher (2007) Electron correlation encoded in this (approximate) wave function can be efficiently dissected in terms of orbital entanglement measuresBoguslawski et al. (2012). Hence an approximate DMRG calculation for all valence orbitals (including also double shells such as $4d$ orbitals for $3d$ transition metal atoms) delivers approximate, but qualitatively correct one- and two-orbital reduced density matrices, from which the single orbital entropies and the mutual informationLegeza and Sólyom (2003); Rissler, Noack, and White (2006); Legeza and Sólyom (2006) are evaluated. Based the single-orbital entropies, autoCAS defines the orbitals for an active space in a following fully converged CASSCF or DMRG calculation, optionally with subsequent dynamic correlation treatment. Even though the autoCAS algorithm is computationally more expensive than other active space selection algorithms, it is based on a multi-configurational description of the wave function and it is therefore not biased toward a single configuration, by contrast to other approaches. Moreover, since the approximate initial DMRG calculation relies on a rather small bond dimension and only a few sweeps to be efficient, the final fully converged calculation allows one to assess whether the initial settings were reliable enough for the active space selection. Should a discrepancy be detected because of truly strong correlations in the electronic structure, then one can easily restart the whole procedure with an increased bond dimension.
@@ II.7.3 API
The Molecule class stores all molecule-specific values, such as spin and charge. It can be initialized through a XYZ file or a list of atoms (strings with element labels). The Autocas class is responsible for the handling of active spaces, either the valence space (number of occupied and virtual orbitals around the Fermi vacuum) or the selected active space based on single-orbital entropies. These functions return 2 lists. One list with the orbital indices and one list storing an occupation number for each orbital with a possible value of 2, 1 or 0, representing a doubly occupied, singly occupied or virtual orbital, respectively. Hence, the total number of electrons in an active space can be determined by summation of the entries in the occupation number list. Even though every interface in autoCAS inherits from a base class, customized workflows only need to provide a list with single-orbital entropies in the same order as the index and occupation lists.
@@ II.8.1 Field of application
II.8.2 I/O

### v3

@@ II.5.1 Field of application
If a system of interest is too large for a full quantum mechanical description, a quantum–classical hybrid model must be constructed. Such models generally describe a small subsystem quantum mechanically (QM) and the remainder with a molecular mechanical (MM) model which parametrizes inter- and intramolecular interactions based on experimental or calculated reference data. The module SCINE Swoose provides such MM models, the framework to combine its models with any existing electronic structure theory to construct a QM/MM hybrid model, and workflows specialized for nanoscopic systems. Generally, MM models can be difficult to construct due to the structural complexity of nanoscopic systems and the necessity of MM models to define bonds and charges a priori. Swoose overcomes this hurdle by providing a general workflow for structure preparationCsizi and Reiher (2023) and a system-focused atomistic model (SFAM)Brunken and Reiher (2020) that can be generated for each system of interest with automated QM reference data generation by fragmenting the system, carrying out numerous structure optimizations in parallel, and fitting force constants with a partial Hessian procedure. The generated structural and MM model can then be leveraged easily in automated proceduresCsizi, Steiner, and Reiher (2023). Reaction mechanisms can be studied with hybrid models, for which the QM region can be automatically determined by another algorithm implemented in SwooseBrunken and Reiher (2021). It generates multiple candidates for the QM region around an atom alongside multiple large reference systems, calculates the first-order partial derivatives for each nucleus, i.e., the atomic forces, and compares them against those in the reference systems. The algorithm assumes that small errors in atomic forces are a good descriptor together with energies as shown in Ref. 38. It then selects the smallest QM system that features errors below a given threshold. Compared to other approaches to determine a QM regionSumner, Söderhjelm, and Ryde (2013); Liao and Thiel (2013); Kulik et al. (2016); Karelina and Kulik (2017); Brandt and Jacob (2023a, b), to determine protonation statesDolinsky et al. (2004, 2007); O’Boyle et al. (2011); Olsson et al. (2011); Anandakrishnan, Aguilar, and Onufriev (2012); Sastry et al. (2013); Riojas and Wilson (2014); Bochevarov et al. (2016); Eastman et al. (2017); Reis et al. (2020), and to fit a system specific modelGrimme (2014); Vanduyfhuys et al. (2015); Welsh and Allison (2019); Behler and Csányi (2021); Friederich et al. (2021); Unke et al. (2021); Deringer et al. (2021); Musil et al. (2021), Swoose exploits the straightforward parallelization of multiple calculations within the SCINE framework and provides a complete pipeline from experimental data to systematic reaction explorations with QM/MM hybrid models, including real-time calculationsCsizi, Steiner, and Reiher (2023).
II.5.2 Scientific background
Swoose provides two molecular mechanics models, GAFFWang et al. (2004) and our system-focused atomistic model SFAM. The latter adopts the following (standard) functional form:
$\begin{split}E_{\mathrm{MM}}&=E_{\mathrm{cov}}+E_{\mathrm{noncov}}=E_{r}+E_{{\alpha}}+E_{{\psi}}+E_{{\phi}}+E_{\mathrm{estat}}\\ &+E_{\mathrm{disp}}+E_{\mathrm{rep}}+E_{\mathrm{hb}}.\end{split}$ (11)
The potential energy contributions from bonds ($E_{r}$) and angles ($E_{{\alpha}}$) are taken to be harmonic potentials. The dihedral angle terms ($E_{{\psi}}$) are modeled by cosine functions. To describe accurately out-of-plane torsions of planar fragments and to account for inversion barriers of trigonal planar fragments, SFAM incorporates an improper dihedral term ($E_{{\phi}}$), which takes a harmonic form for small out-of-plane angles and a double-well potential for large equilibrium angles. The electrostatic term ($E_{\mathrm{estat}}$) is defined as a pairwise sum of Coulomb interactions between atom-centered partial charges. For repulsive ($E_{\mathrm{rep}}$) and dispersive ($E_{\mathrm{disp}}$) interactions, we adopt the functional form of Grimme’s QMDFF force fieldGrimme (2014), which consists of an exponential term for repulsions alongside the D3 semiclassical dispersion correctionsGrimme et al. (2010). Also the SFAM-hydrogen bonding potentials ($E_{\mathrm{hb}}$) for atom triples (hydrogen donor, hydrogen acceptor, hydrogen atom) are inspired by QMDFF.
The amount of required reference data to derive all parameters in Eq. (11) for any given structure is rather limited and only comprises the equilibrium structure of the molecule, the corresponding Hessian matrix, the atomic partial charges, and the covalent bond orders. Divide-and-conquer fragmentation schemes allow us to provide scalable variants of the parametrization routine applicable to the nanoscopic scale. For details on the functional form and parametrization procedures, we refer to Ref. 37.
After successful parametrization, SFAM can be employed as a classical force field in hybrid QM/MM calculations. As SFAM parameters are always generated for the entire molecular system, the QM region can be dynamically redefined at any time. Furthermore, SFAM can be locally reparametrized if the bonding pattern in the QM region changes through a chemical reaction, but the corresponding atoms are subsequently transferred to the MM region. In the QM/SFAM model, both mechanical embedding (ME) and electrostatic embedding (EE) are supported. Although the quantum mechanical treatment of interaction energies in the EE scheme is more accurate, the applicability of EE depends on the chosen QM model and on whether the QM Hamiltonian can be equipped with one-electron terms for the charged MM atoms. Within SCINE, xTB and the external programs ORCA and Turbomole can be employed for EE-QMM/MM calculations. For ME, the QM/SFAM energy expression of the nanoscopic system composed of an environment region $\mathcal{E}$ and a quantum region $\mathcal{Q}$ is defined as
$E_{\mathrm{QM/SFAM}}^{\mathrm{ME}}=E_{\mathrm{QM,ME}}^{\mathcal{Q}}+E_{\mathrm{SFAM}}^{\mathcal{Q}+\mathcal{E}}-E_{\mathrm{SFAM}}^{\mathcal{Q}}$ (12)
The latter part subtracts all contributions that are covered by the QM calculation to avoid double counting. For EE, the QM/SFAM energy expression reads
$E_{\mathrm{QM/SFAM}}^{\mathrm{EE}}=E_{\mathrm{QM/SFAM}}^{\mathrm{ME}}+E_{\mathrm{QM,estat}}^{\mathcal{Q}}-\sum_{\begin{subarray}{c}A\in\mathcal{Q}\\ B\in\mathcal{E}\end{subarray}}E_{\mathrm{estat}}^{AB},$ (13)
with the QM electronic energy obtained with electrostatic embedding. To account for the unfavorable scaling of the error in the total energy for large molecular systems, we proposed a reduced QM/MM energy:
$E_{\mathrm{QM/SFAM}}^{\mathrm{red}}=E_{QM}^{\mathcal{Q}}-E^{\mathcal{Q}-\mathcal{E}},$ (14)
where $E^{\mathcal{Q}-\mathcal{E}}$ is the interaction energy of either ME or EE type that is captured by the QM method. This formalism neglects covalent SFAM contributions and noncovalent interactions between environment atoms, and therefore, it compensates the potentially large uncertainties induced by large environments $\mathcal{E}$.
QM/SFAM also provides a universal procedure to select an accurate QM region as the reactive center. To this end, candidate QM regions around a central atom are generated leveraging the SFAM fragmentation approach coupled to a stochastic element (with bonds being cut with some probability $p$). The candidate regions are selected within an upper and lower boundary in terms of atom count. These parameters allow to flexibly define the QM region size and variation thereof. For that purpose, the mean absolute error of the force components on atoms that are closer to the selected central atom than a give threshold $r_{\text{repr}}$ in the candidate regions with respect to the larger reference regions are calculated. The mean absolute error of one atom $k$ with regard to the reference is defined as
$e_{k}^{m}=\frac{1}{3}\Big{(}|f_{x,\text{ref},k}-f_{x,m,k}|+|f_{y,\text{ref},k}-f_{y,m,k}|+|f_{z,\text{ref},k}-f_{z,m,k}|\Big{)},$ (15)
and the mean of all representative atoms as
$e_{\text{mean}}^{k}=\frac{1}{N_{\text{repr}}}\sum_{k=1}^{N_{\text{repr}}}e_{k}^{m}.$ (16)
In Eq. (15), $\mathbf{f}_{\text{ref},k}=(f_{x,\text{ref},k},f_{y,\text{ref},k},f_{z,\text{ref},k})^{T}$ is the reference force vector of atom $k$ and can be calculated with any QM method. The generated candidate regions can also be scored by the number of cleaved bonds at the QM/MM boundary (and hence, the number of inserted link atoms), and by the overall symmetry of the QM region.
To employ QM/SFAM for optimization routines provided within the SCINE infrastructure, we provide dedicated QM/MM optimizers that find local minima and QM/MM transition states on a highly complex potential energy surface. Both optimizers rely on a microiterative structure: For structure optimizations, all Cartesian coordinates of the quantum region and those of the environment that are bound to the quantum region are constrained, while all remaining MM degrees of freedom are relaxed until convergence (or until a maximum number of iterations is reached). After that, the complete system is optimized to a local minimum applying the full analytical QM/SFAM gradients.
For transition state optimizations, we exploit a partial Hessian scheme to optimize all atoms in the QM region while, at the same time, converging the atoms in the MM region to a local minimum. This approach avoids the calculation of MM Hessians. All MM atom positions are iteratively updated as point charges in the QM Hamiltonian after convergence of the MM optimization. Then, the QM region is optimized by applying a Bofill update (eigenvector following), and the QM positions are updated in the full model after the maximum number of iterations. As the subsystems are optimized with individual optimizers, convergence is achieved when the QM optimization converges twice (before and after an MM optimization) in order to guarantee self-consistency.
II.5.3 I/O
@@
The algorithms for structure preparation, model generation, and quantum region selection can be carried out with Swoose with an input file following the YAML syntax, which allows one to specify all required settings in a straightforward way. Additionally, Swoose can carry out single-point energy calculations, structure optimizations, and molecular dynamics simulations directly on the command line.
II.5.4 API
@@ II.7.1 Field of application
Highly accurate electronic structure calculations require one to take electron correlation into consideration. Even though closed-shell molecules can be routinely and reliably calculated with black-box coupled-cluster modelsBartlett (2024), strongly correlated systems require an active orbital space for a multi-configurational description with subsequent treatment of the dynamic correlation. Since the manual selection of an active space requires chemical intuition and experience, this step is often time consuming, difficult to reproduce, and hardly possible for non-prototypical complex system such as the one in Ref. 153. Especially for methods which can handle large active spaces of more than 100 orbitals, like DMRGWhite (1992, 1993); Baiardi and Reiher (2020) and full configuration interaction quantum Monte CarloBooth, Thom, and Alavi (2009); Cleland, Booth, and Alavi (2010), a manual selection is neither feasible nor reliable. Although many approachesPulay and Hamilton (1988); Bofill and Pulay (1989); Tishchenko, Zheng, and Truhlar (2008); Bao et al. (2016, 2017); Sayfutyarova et al. (2017); Bao et al. (2018); Khedkar and Roemelt (2019); Sayfutyarova and Hammes-Schiffer (2019); Jeong et al. (2020); Li, Gagliardi, and Truhlar (2020); Gieseking (2021); Golub et al. (2021); Khedkar and Roemelt (2021); Lei, Suo, and Liu (2021); Levine et al. (2021); Oakley, Gagliardi, and Truhlar (2021); King and Gagliardi (2021); Weser et al. (2022); Casanova (2022); Cheng, Xie, and Ma (2022); King et al. (2022); Kaufold et al. (2023); Golub et al. (2023), have been devised to address this problem (at least in parts), we have demonstrated in 2016 a viable, reliable, and truly fully automated selection approach based on quantum information measures, the autoCAS approachStein, von Burg, and Reiher (2016); Stein and Reiher (2016, 2017, 2019); Bensberg and Reiher (2023c), a development, which has not been paralleled so far.
@@
The autoCAS algorithmStein and Reiher (2016, 2019) exploits the fact that DMRG can converge a qualitatively correct wave function with less effort than what would be required for a converged energyMoritz and Reiher (2007). Electron correlation encoded in this (approximate) wave function can be efficiently dissected in terms of orbital entanglement measuresBoguslawski et al. (2012). Hence an approximate DMRG calculation for all valence orbitals (including also double shells such as $4d$ orbitals for $3d$ transition metal atoms) delivers approximate, but qualitatively correct one- and two-orbital reduced density matrices, from which the single orbital entropies and the mutual informationLegeza and Sólyom (2003); Rissler, Noack, and White (2006); Legeza and Sólyom (2006) are evaluated. Based on the single-orbital entropies, autoCAS defines the orbitals for an active space in a following fully converged CASSCF or DMRG calculation, optionally with subsequent dynamic correlation treatment. Even though the autoCAS algorithm is computationally more expensive than other active space selection algorithms, it is based on a multi-configurational description of the wave function and it is therefore not biased toward a single configuration, by contrast to other approaches. Moreover, since the approximate initial DMRG calculation relies on a rather small bond dimension and only a few sweeps to be efficient, the final fully converged calculation allows one to assess whether the initial settings were reliable enough for the active space selection. Should a discrepancy be detected because of truly strong correlations in the electronic structure, then one can easily restart the whole procedure with an increased bond dimension.
@@ II.7.3 API
The Molecule class stores all molecule-specific values, such as spin and charge. It can be initialized through a XYZ file or a list of atoms (strings with element labels). The Autocas class is responsible for the handling of active spaces, either the valence space (number of occupied and virtual orbitals around the Fermi vacuum) or the selected active space based on single-orbital entropies. These functions return 2 lists: one list with the orbital indices and one list storing an occupation number for each orbital with a possible value of 2, 1 or 0, representing a doubly occupied, singly occupied or virtual orbital, respectively. Hence, the total number of electrons in an active space can be determined by summation of the entries in the occupation number list. Even though every interface in autoCAS inherits from a base class, customized workflows only need to provide a list with single-orbital entropies in the same order as the index and occupation lists.
@@ II.8.1 Field of application
II.8.2 Scientific background
Some of the algorithms accessible through ReaDuct are well-established in the literature, while others were specifically designed within the SCINE project. One example of the latter are the so-called Newton trajectory tasks, originally introduced in Ref. 32 and later improved upon in Ref. 28. Their development was inspired by the work of Quapp and Bofill on Newton trajectoriesQuapp and Bofill (2020) (hence the name) and their analysis of Maeda’s artificial force induced reaction (AFIR) approachMaeda and Morokuma (2010, 2011); Maeda, Taketsugu, and Morokuma (2014). The Newton Trajectory task 2 (NT2) algorithm has become our workhorse for single-ended transition state searches, which is why we describe the underlying algorithm in the following.
Starting from a reactant’s minimum energy structure (or a reactive complex composed of several reactants) NT2 scans aim to locate transition state guesses in a given direction. This direction is specified by a list of so-called reactive atom pairs, referred to as the reactive coordinate. Each pair is either set to be pushed together (associative) or to be pulled apart (dissociative). For an illustration see Fig. 5.
Figure 5: Reactive coordinates in NT2 tasks. Reactive pairs (two circles of same color) can be associative (green arrows) or dissociative (orange arrows). Reactive coordinates may include associative or dissociative reactive pairs only (left and center) or both types of pairs simultaneously (right). Figure adapted from Ref. 196.
Note that while the selection of these pairs can be motivated by the expectation of certain bonds to be formed or broken, the algorithm neither strictly enforces specific bond modifications nor does it limit changes in the molecular graph to the selected atom pairs.
An NT2 scan is a steepest descent (SD) structure optimization under the application of a modified gradient, such that the system is pushed energetically uphill into the direction of interest. This modified gradient $\mathbf{g}^{\text{NT2}}$ is obtained from the original electronic energy gradient $\mathbf{g}$ as follows: While the gradient components acting on atoms that are not included in the reactive coordinate remain unchanged, the gradient of each reactive atom $k$ is orthogonalized with respect to the vectors $\mathbf{R}_{a_{i}b_{i}}$ connecting the reactive atom pairs $\mathbf{P}^{\text{react}}_{i}=(a_{i},b_{i})$ to which $k$ belongs:
$\mathbf{g}_{k}\rightarrow\textbf{g}_{k}^{\perp},\quad\text{with}\quad\textbf{g}_{k}^{\perp}\perp\textbf{R}_{a_{i}b_{i}}\quad\forall\textbf{P}_{i}^{\text{react}}=(a_{i},b_{i})\ni k$ (17)
Subsequently, artificial force components are included,
$\mathbf{g}_{k}^{\text{NT2}}=\textbf{g}_{k}^{\perp}-0.5\frac{\alpha_{\text{NT2}}}{\text{max}_{\forall(a_{i},b_{i})\ni k}||\mathbf{X}_{a_{i}b_{i}}||}\cdot\left(\sum_{i}\mathbf{X}_{kb_{i}}-\sum_{i}\mathbf{X}_{a_{i}k}\right),$ (18)
where $\alpha_{\text{NT2}}$ is a positive force constant. The two sums run over all reactive pairs with atom $k$ being their first or second element. The calculation of $\mathbf{X}_{a_{i}b_{i}}$ depends on whether the pair $(a_{i},b_{i})$ was set to be pushed together or to be pulled apart. For the associative case $\mathbf{X}_{a_{i}b_{i}}$ is
$\mathbf{X}_{a_{i}b_{i}}^{\text{associative}}=\left(||\mathbf{R}_{a_{i}b_{i}}||-\left(r_{a_{i}}^{\text{cov}}+r_{b_{i}}^{\text{cov}}\right)\right)\cdot\frac{\mathbf{R}_{a_{i}b_{i}}}{||\mathbf{R}_{a_{i}b_{i}}||}$ (19)
with $||\mathbf{R}_{a_{i}b_{i}}||$ denoting the distance between atoms $a_{i}$ and $b_{i}$ and $r_{a_{i}}^{\text{cov}}$ and $r_{b_{i}}^{\text{cov}}$ their covalent radii, respectively. Hence, the prefactor $||\mathbf{R}_{a_{i}b_{i}}||-\left(r_{a_{i}}^{\text{cov}}+r_{b_{i}}^{\text{cov}}\right)$ scales the force component according to the difference between the actual interatomic distance and an idealized bond distance. This facilitates concerted bond formations when multiple associative pairs are considered simultaneously. For pairs that are to be pulled apart $\mathbf{X}_{a_{i}b_{i}}$ is the reversed, normalized intra-pair connection vector scaled with a factor $\lambda$,
$\mathbf{X}_{a_{i}b_{i}}^{\text{dissociative}}=-\lambda\cdot\frac{\mathbf{R}_{a_{i}b_{i}}}{||\mathbf{R}_{a_{i}b_{i}}||}$ (20)
Adjusting $\lambda$ allows one to control the magnitude of associative vs. dissociative components of the overall artificial force.
Each SD step is followed by BFGS optimization cycles during which the positions of all atoms that are part of the reactive coordinate are frozen, while the remainder of the structure relaxes, i.e., the remainder adapts to the induced changes. One SD step and the subsequent BFGS relaxation steps are referred to as a macrocycle.
The scan is stopped (i) upon reaching a user-specified number of gradient calculations or (ii) if the distance between reactive pairs being pushed together is smaller than the sum of their covalent radii times 0.9 or the bond order between them exceeds 0.75, and the bond order between atom pairs that are pulled apart is below 0.15. Furthermore, the scan finishes when a gradient calculation fails. A transition state guess structure is extracted from the trajectory obtained from the macrocycles of the scan (i.e., post-BFGS relaxation) based on changes in the bonding pattern within the reactive pairs and the electronic energy curve (for details see Ref. 28). The obtained guess structure can then be subjected to a transition state optimization as outlined in section II.8.1.
II.8.3 I/O